# Uniform Feasibility For Smoothed Backup Control Barrier Functions

### v1

@@ V Conclusion
We have established uniform feasibility guarantees for CBF-based controllers such as safety filters using converse safety theorems when the safe set is constructed using smoothened minima of $C^{1}$ functions (via soft-min). For compact safe sets, an explicit parameter is designated for certifying the soft-min function as a CBF, which implies the feasibility of the CBF-based safety constraint. For noncompact sets, a set of tail conditions was shown to yield an extended CBF (a more relaxed formulation of CBF) with the same feasibility results. Projecting theoretical findings onto the backup CBFs, we show that strict safety of a compact backup set under a backup controller, combined with a strict safety condition along the relevant boundary of the safe set, is sufficient for feasibility in nonsmooth backup CBF formulation.
Future work will focus on relaxing assumptions of compact backup set and strict positiveness on the backup-reachable boundary of the safe set. Another direction is to investigate regularity (local Lipshitzness) and robustness properties of the smoothed BCBF framework, including sensitivity to model mismatch and sampling/discretization effects.

### v2

@@ VI Conclusion
We have established uniform feasibility guarantees for CBF-based controllers such as safety filters using converse safety theorems when the safe set is constructed using smoothed minima of $C^{1}$ functions (via soft-min). For compact safe sets, an explicit parameter is designated for certifying the soft-min function as a CBF, which implies the feasibility of the CBF-based safety constraint. For noncompact sets, a set of tail conditions was shown to yield an extended CBF (a more relaxed formulation of CBF) with the same feasibility results. Projecting theoretical findings onto the BCBFs, we show a set of sufficient conditions for feasibility in BCBF formulation. Future work will focus on relaxing assumptions of compact backup set on the backup-reachable boundary of the safe set. Investigating the time-varying extension is another direction for future work.